# Who Made This? Fake Detection and Source Attribution with Diffusion Features

### v1

@@ 6 Conclusions
In this work, we presented FRIDA, a training-free framework that repurposes internal activations of pre-trained diffusion models for deepfake detection and source generator attribution. FRIDA operates entirely in the diffusion feature space, leveraging the latent representation of a Stable Diffusion U-Net as discriminative descriptors of image authenticity. A simple k-nearest-neighbor classifier applied to these features achieves state-of-the-art detection performance and generalizes effectively across unseen generators, while a compact neural model trained on the same representations enables accurate source attribution. Our analysis shows that diffusion features, although optimized for image generation, are surprisingly effective for fake image detection and source model attribution. This finding suggests that diffusion features can serve as a universal and interpretable basis for synthetic image forensics, bridging generative modeling and authenticity analysis.

### v2

@@ 6 Conclusions
In this work, we introduced FRIDA, a framework that leverages the internal representation of a pre-trained diffusion model for fake image detection and source model attribution. For fake image detection, FRIDA entirely operates within the diffusion feature space, exploiting latent representations from the Stable Diffusion U-Net as compact and discriminative descriptors of image authenticity. Through extensive evaluation, we demonstrated that a simple $k$-nearest neighbour classifier applied to these features achieves state-of-the-art detection performance with a limited support set, generalizing effectively to unseen generators without any retraining. For source model attribution, a compact neural classifier trained on the same features accurately identifies the generative source, confirming that diffusion features encode model-specific fingerprints. Furthermore, we show that diffusion-based features are more discriminative than CLIP and DINO features for fake image detection. Through extensive testing, we also demonstrate that our method maintains high performance under various image distortions and effectively generalizes to out-of-distribution generators, including state-of-the-art models like Flux and Qwen. These findings may position diffusion representations as a reliable foundation for AI-generated image forensics.